OpenStudy (anonymous):

solve by the elimination method,7r -5s=27, 5r+7s=51

OpenStudy (anonymous):

Equation 1: 7r-5s=27 Equation 2: 5r+7s=51 Multiply both to LCM (least common multiple, in this case 35) 35r-25s=135 35r+49s=357 Equation 2 minus equation 1 = 24s=222 You can now find s, then use the value for s to solve for r.